# Bring Your Dreams to Life: Continual Text-to-Video Customization

### v1

@@ Abstract
Customized text-to-video generation (CTVG) has recently witnessed great progress in generating tailored videos from user-specific text. However, most CTVG methods assume that personalized concepts remain static and do not expand incrementally over time. Additionally, they struggle with forgetting and concept neglect when continuously learning new concepts, including subjects and motions. To resolve the above challenges, we develop a novel Continual Customized Video Diffusion (CCVD) model, which can continuously learn new concepts to generate videos across various text-to-video generation tasks by tackling forgetting and concept neglect. To address catastrophic forgetting, we introduce a concept-specific attribute retention module and a task-aware concept aggregation strategy. They can capture the unique characteristics and identities of old concepts during training, while combining all subject and motion adapters of old concepts based on their relevance during testing. Besides, to tackle concept neglect, we develop a controllable conditional synthesis to enhance regional features and align video contexts with user conditions, by incorporating layer-specific region attention-guided noise estimation. Extensive experimental comparisons demonstrate that our CCVD outperforms existing CTVG models. The code is available at https://github.com/JiahuaDong/CCVD.
Introduction
@@
Figure 1: Results of our model in continually learning new personalized concepts, including subjects and motions, for customized text-to-video generation. Compared with L2DM, our model achieves better performance to address the CTVC problem.
@@
To the CTVC problem, we propose a novel Continual Customized Video Diffusion (CCVD) model, which is the first work to tackle catastrophic forgetting and concept neglect across a variety of CTVG tasks. First, to overcome catastrophic forgetting of previously learned personalized concepts, we develop a concept-specific attribute retention (CAR) module for training and a task-aware concept aggregation (TCR) strategy for testing. The CAR module incorporates layer-specific concept tokens and a concept orthogonal loss to explore the unique characteristics and identity of personalized concepts, while the TCR module merges all subject and motion adapters of old concepts, weighted by their relevance to CTVG tasks. Second, we design a controllable conditional synthesis module to address concept neglect in the multi-concept video customization task. It employs layer-specific region attention to enhance the representational capacity of latent regional features and utilizes attention-guided noise estimation to control the context of personalized subjects and motions within the generated video, aligning with user-specified conditions. Experiments demonstrate the effectiveness of our CCVD. The key contributions are as follows:
@@
Preliminary and Problem Definition
@@
Figure 3: Comparison results of single-concept video customization under the CTVC setting.
@@
Figure 4: Comparison results of multi-concept video customization under the CTVC setting.
Figure 5: Comparisons of style transfer under CTVC setting.
Experiments
@@
Figure 6: Comparisons of video editing under CTVC setting.
Table 1: Results of single-concept video customization.
@@ Methods
Setups: For fair comparisons, our model and all comparative methods utilize DreamVideo (10656166) as the baseline backbone. We use the Adam optimizer for training, with an initial learning rate of $1.0\times 10^{-4}$ for textual embeddings and $1.0\times 10^{-5}$ for the 3D UNet. We set $\lambda=0.1$ in Eq. (2), and $n_{v}=32$ to synthesize customized videos. Besides, we perform comprehensive qualitative and quantitative evaluations across various generation tasks including single/multi-concept video customization, style transfer and video editing. For quantitative comparisons, we follow (10656166) to utilize CLIPT, CLIPI, DINOI, TCons as evaluation metrics. Moreover, we use the forgetting of CLIPT, CLIPI, and DINOI (denoted as F-CLIPT, F-CLIPI, and F-DINOI) (sun2024create) to assess CTVC performance.
Qualitative Comparisons
To demonstrate the superiority of our model, as in Fig. 36, we conduct qualitative comparisons on single/multi-concept video customization, style transfer and video editing. 1) In Fig. 3, our model surpasses other methods in single-concept video customization, since it can mitigate forgetting by incorporating concept-specific attribute retention module to explore unique concept identity. 2) In Fig. 4, all comparative methods fail to generate multiple concepts, while our model can leverage layer-specific region attention and attention-guided noise estimation to synthesize videos with multiple user-specified concepts. 3) CCVD achieves superior performance in video style transfer (Fig. 5), as our task-aware concept aggregation module can better maintain style integrity. 4) In Fig. 6, we integrate Video-P2P (10656631) as a plugin for editing. The large improvement of CCVD compared to other methods is attributed to the concept-specific attribute retention module, which can preserve unique concept.
Table 2: Results (CLIPI) of single-concept customization.
@@
Table 3: Ablation of single-concept video customization.
@@
To verify the efficacy of all proposed modules, we conduct comprehensive ablation experiments in terms of layer-specific concept tokens (LCT), concept orthogonal loss (COL), task-aware concept aggregation (TCA), layer-specific region attention (LRA) and attention-guided noise estimation (ANE). As shown in Tab. 3, compared to “Baseline”, our model achieves significant performance improvements across all evaluation metrics after incorporating the proposed LCT, COL, and TCA modules for single-concept video customization. It illustrates that our model effectively mitigates forgetting by capturing distinctive attributes of old concepts while dynamically aggregating subject and motion adapters. It confirms their effectiveness in mitigating concept neglect. For more detailed analysis and visualization, please refer to our supplementary materials and anonymous project page.
Conclusion

### v2

@@ Abstract
Customized text-to-video generation (CTVG) has recently witnessed great progress in generating tailored videos from user-specific text. However, most CTVG methods assume that personalized concepts remain static and do not expand incrementally over time. Additionally, they struggle with forgetting and concept neglect when continuously learning new concepts, including subjects and motions. To resolve the above challenges, we develop a novel Continual Customized Video Diffusion (CCVD) model, which can continuously learn new concepts to generate videos across various text-to-video generation tasks by tackling forgetting and concept neglect. To address catastrophic forgetting, we introduce a concept-specific attribute retention module and a task-aware concept aggregation strategy. They can capture the unique characteristics and identities of old concepts during training, while combining all subject and motion adapters of old concepts based on their relevance during testing. Besides, to tackle concept neglect, we develop a controllable conditional synthesis to enhance regional features and align video contexts with user conditions, by incorporating layer-specific region attention-guided noise estimation. Extensive experimental comparisons demonstrate that our CCVD outperforms existing CTVG baselines on both the DreamVideo and Wan 2.1 backbones. The code and models are available at https://github.com/JiahuaDong/CCVD.
1 Introduction
@@
Figure 1: Results of our model in continually learning new personalized concepts, including subjects and motions, under the CTVC setting. Compared with L2DM, our model achieves better performance on both the DreamVideo and Wan 2.1 backbones.
@@
To the CTVC problem, we propose a novel Continual Customized Video Diffusion (CCVD) model, which is the first work to tackle catastrophic forgetting and concept neglect across a variety of CTVG tasks. First, to overcome catastrophic forgetting of previously learned personalized concepts, we develop a concept-specific attribute retention (CAR) module for training and a task-aware concept aggregation (TCR) strategy for testing. The CAR module incorporates layer-specific concept tokens and a concept orthogonal loss to explore the unique characteristics and identity of personalized concepts, while the TCR module merges all subject and motion adapters of old concepts, weighted by their relevance to CTVG tasks. Second, we design a controllable conditional synthesis module to address concept neglect in the multi-concept video customization task. It employs layer-specific region attention to enhance the representational capacity of latent regional features and utilizes attention-guided noise estimation to control the context of personalized subjects and motions within the generated video, aligning with user-specified conditions. Experiments demonstrate the effectiveness of our CCVD model on both the DreamVideo and Wan 2.1 backbones. The key contributions are as follows:
@@ 2 Related Work
3 Preliminary and Problem Definition
@@ 4 The Proposed Model
Figure 3: Comparison results of single-concept video customization under the CTVC setting when the backbone is DreamVideo.
@@
Figure 4: Comparison results of multi-concept video customization under the CTVC setting when the backbone is DreamVideo.
Figure 5: Comparisons of style transfer under CTVC setting when the backbone is DreamVideo.
5 Experiments
@@
Figure 6: Comparisons of video editing under CTVC setting when the backbone is DreamVideo.
Figure 7: Comparison results of single-concept video customization under the CTVC setting when the backbone is Wan 2.1.
Figure 8: Comparison results of multi-concept video customization under the CTVC setting when the backbone is Wan 2.1.
Table 1: Results of single-concept video customization when the backbone is DreamVideo.
@@ Methods
Table 2: Results (CLIPI) of single-concept customization when the backbone is DreamVideo.
@@
Setups: For fair comparisons, our model and all comparative methods utilize DreamVideo (10656166) and Wan 2.1 (wan2025wanopenadva) as the baseline backbones. For DreamVideo (10656166), we use the Adam optimizer for training, with an initial learning rate of $1.0\times 10^{-4}$ for textual embeddings and $1.0\times 10^{-5}$ for the 3D UNet. For Wan 2.1 (wan2025wanopenadva), the initial learning rate for both the textual embeddings and the diffusion transformer backbone is set to $1.0\times 10^{-4}$. Besides, we set $\lambda=0.1$ in Eq. (2), and $n_{v}=32$ to synthesize customized videos. Besides, we perform comprehensive qualitative and quantitative evaluations across various generation tasks including single/multi-concept video customization, style transfer and video editing. For quantitative comparisons, we follow (10656166) to utilize CLIPT, CLIPI, DINOI, TCons as evaluation metrics. Moreover, we use the forgetting of CLIPT, CLIPI, and DINOI (denoted as F-CLIPT, F-CLIPI, and F-DINOI) (sun2024create) to assess CTVC performance.
Qualitative Comparisons
To demonstrate the superiority of our CCVD model, as in Fig. 38, we conduct qualitative comparisons on single/multi-concept video customization, style transfer and video editing. 1) In Figs. 3 and 7, our model surpasses other methods in single-concept video customization, since it can mitigate forgetting by incorporating concept-specific attribute retention module to explore unique concept identity. 2) In Fig. 4 and 8, all comparative methods fail to generate multiple concepts, while our model can leverage layer-specific region attention and attention-guided noise estimation to synthesize videos with multiple user-specified concepts. 3) CCVD achieves superior performance in video style transfer (Fig. 5), as our task-aware concept aggregation module can better maintain style integrity. 4) In Fig. 6, we integrate Video-P2P (10656631) as a plugin for editing. The large improvement of CCVD compared to other methods is attributed to the concept-specific attribute retention module, which can preserve unique concept.
@@
Table 3: Ablation of single-concept video customization when the backbone is DreamVideo.
@@
To verify the efficacy of all proposed modules, we conduct comprehensive ablation experiments in terms of layer-specific concept tokens (LCT), concept orthogonal loss (COL), task-aware concept aggregation (TCA), layer-specific region attention (LRA) and attention-guided noise estimation (ANE). As shown in Tab. 3, compared to “Baseline”, our model achieves significant performance improvements across all evaluation metrics after incorporating the proposed LCT, COL, and TCA modules for single-concept video customization. It illustrates that our model effectively mitigates forgetting by capturing distinctive attributes of old concepts while dynamically aggregating subject and motion adapters. As presented in Fig. 9, the performance of our model in multi-concept video customization significantly declines when the LRA and ANE modules are removed. This confirms their effectiveness in mitigating the issue of concept neglect.
Figure 9: Ablation of multi-concept video customization when the backbone is DreamVideo.
6 Conclusion
@@ Acknowledgments
Appendix A Supplementary Material
Implementation Details
Dataset: Inspired by (1011453687945; 10656166), we introduce a new continual video customization (CVC) dataset to evaluate the effectiveness of our model under the CTVC setting. As shown in Fig. 10, there are 35 sequential text-to-video customization tasks in the CVC dataset, including 25 personalized subjects and 10 motion patterns. Among the 25 personalized subjects, 9 are pets, 5 are styles, and the remaining 11 are objects. To satisfy the practical requirements of the CTVC setting, during training, we set one text-video pair for tasks involving motions and $3\sim 5$ image-text pairs for tasks involving subjects. More importantly, we include several semantically similar personalized concepts (e.g., seven dog concepts and two cat concepts) to make the CVC dataset more challenging.
Figure 10: Visualization of 35 consecutive text-to-video customization tasks in the CVC dataset.
Figure 11: Text prompts used in the paper for quantitative comparisons.
Setups: For fair comparisons, our proposed model and all comparative methods utilize DreamVideo (10656166) as the baseline backbone. For DreamVideo (10656166), we use the Adam optimizer for training, with an initial learning rate of $1.0\times 10^{-4}$ for textual embeddings and $1.0\times 10^{-5}$ for the 3D UNet. For Wan 2.1 (wan2025wanopenadva), the initial learning rate for both the textual embeddings and the diffusion transformer backbone is set to $1.0\times 10^{-4}$. In this paper, we set $\lambda=0.1$ in Eq. (2). Furthermore, we set the training steps to $1,000$ for subject and motion adapters and $3,000$ for concept tokens. During the inference phase, we follow DreamVideo (10656166) to employ the DDIM sampler (song2021denoising) with 50 steps and apply classifier-free guidance (ho2021classifierfree) across all baseline models. Moreover, videos are generated with a spatial resolution of $256\times 256$, consisting of $32$ frames at a rate of 8 frames per second (fps). All comparison experiments are performed on a single NVIDIA 3090 GPU.
Table 4: Results (CLIPI) of single-concept customization when the backbone is DreamVideo.
Methods
V1-5 V6-10 V11-15 V15-20 V21-25 V25-30 V31-35  Avg.
Finetuning 21.9 23.1 29.8 22.9 23.0 22.6 23.0 22.9
EWC 30.0 28.1 28.8 28.5 27.6 27.6 25.8 28.1
LWF 29.3 24.1 23.7 25.7 24.2 23.1 23.7 24.8
LoRA-M 27.6 25.5 25.2 25.7 25.9 23.7 24.6 25.5
CLoRA 30.3 29.0 29.4 28.8 27.9 28.3 27.0 28.7
L2DM 29.6 28.2 29.2 28.1 27.7 27.9 26.9 28.3
CCVD 31.1 29.7 30.6 30.2 28.9 29.0 29.3 29.9
Evaluation Metrics: We perform comprehensive qualitative and quantitative evaluations across various text-to-video generation tasks including single/multi-concept video customization, style transfer and video editing. As illustrated in Fig. 11, we broadly classify prompts into four groups based on the concepts they encompass. To conduct quantitative comparisons between our model and state-of-the-art (SOTA) methods, following the approach of (1011453687945), we utilize an evaluation prompt and a consistent negative prompt to generate 10 videos. Consequently, with 10 evaluation prompts provided in this study, we produce $100$ videos per concept to assess model performance. Motivated by (10656166), we utilize CLIPT, CLIPI, DINOI, TCons as evaluation metrics for quantitative comparisons. Moreover, we introduce the forgetting of CLIPT, CLIPI, and DINOI (denoted as F-CLIPT, F-CLIPI, and F-DINOI) to assess our model’s performance in addressing catastrophic forgetting of previously learned concepts.
Comparative Methods: To thoroughly showcase the superior performance of our model, we introduce six state-of-the-art (SOTA) comparison methods. They include three continual learning-based approaches (e.g., Finetuning, EWC (kirkpatrick2017overcoming) and LWF (li2017learning)), two continual diffusion models (e.g., CLoRA (smith2023continual) and L2DM (sun2024create)), and one multi-concept composition technology (LoRA-M (zhong2024multi)). Specifically, Finetuning focuses on optimizing subject and motion adapters to sequentially learn multiple personalized concepts. To mitigate catastrophic forgetting, EWC (kirkpatrick2017overcoming) employs an elastic regularizer to constrain changes in network parameters, and LWF (li2017learning) retains old training data to facilitate knowledge distillation for the current task. CLoRA (smith2023continual) introduces a self-regularized low-rank adaptation mechanism to continuously learn new personalized concepts. L2DM (sun2024create) constructs a long-term memory bank to reconstruct past personalized concepts and utilizes knowledge distillation to reduce catastrophic forgetting. LoRA-M (zhong2024multi) combines all subject and motion adapters equally to retrain the latent diffusion model. More importantly, all comparison methods and our proposed model employ DreamVideo (10656166) as the baseline backbone for fair comparisons.
Table 5: Results (DINOI) of single-concept customization when the backbone is DreamVideo.
Methods
V1-5 V6-10 V11-15 V15-20 V21-25 V25-30 V31-35  Avg.
Finetuning 28.1 22.4 31.5 27.3 22.5 24.7 19.3 25.1
EWC 33.2 30.5 35.2 32.0 24.9 25.9 24.6 29.5
LWF 33.9 31.9 36.2 32.8 25.5 25.6 24.6 30.0
LoRA-M 32.9 30.2 35.0 31.9 24.7 24.3 24.0 29.0
CLoRA 34.3 33.1 37.6 32.6 27.2 26.5 25.4 31.0
L2DM 36.2 33.6 35.6 32.5 27.7 26.6 25.9 31.2
CCVD 38.3 34.8 39.1 35.5 28.4 29.7 28.0 33.4
Table 6: Results (F-CLIPI) of single-concept customization when the backbone is DreamVideo.
Methods
V1-5 V6-10 V11-15 V15-20 V21-25 V25-30 V31-35  Avg.
Finetuning 12.7 12.9 14.5 14.9 15.4 15.6 15.5 14.5
EWC 6.2 5.9 7.4 8.1 8.8 9.3 9.4 7.8
LWF 5.3 5.2 5.9 6.8 7.8 8.4 8.4 6.8
LoRA-M 7.1 6.1 7.9 8.6 9.5 9.9 10.0 8.4
CLoRA 3.6 3.5 4.4 5.3 6.4 7.0 7.1 5.3
L2DM 2.6 2.9 3.7 4.0 4.3 4.7 4.8 3.9
CCVD 1.0 1.3 1.9 2.4 2.8 3.3 3.5 2.3
Quantitative Comparisons
To perform quantitative evaluations between our proposed CCVD model and other competing approaches, we adhere to the evaluation framework outlined in (1011453687945). Specifically, we define 10 prompts for each concept and produce 10 videos per prompt, yielding a total of 100 generated videos for performance analysis. As demonstrated in Tabs. 46, our proposed model outperforms competing methods across all evaluation metrics, confirming its superiority in addressing the CTVC problem. This highlights the effectiveness of our model in preserving the distinct identity of each learned concept through the concept-specific attribute retention module. Additionally, the proposed task-aware concept aggregation strategy dynamically combines the subject and motion adapters, effectively mitigating catastrophic forgetting of previously learned concepts during the testing phase.
Figure 12: Comparison results of single-concept video customization under the CTVC setting when the backbone is DreamVideo.
Figure 13: Comparison results of single-concept video customization under the CTVC setting when the backbone is DreamVideo.
Qualitative Comparisons
Single-Concept Video Customization: As presented in Figs. 1213, we perform in-depth qualitative evaluations of single-concept video customization. In Fig. 12, when provided with a text prompt (e.g., “a panda performing V3 in the forest”) for inference, our model accurately adheres to the prompt instructions, producing videos that maintain the distinctive identity of the personalized concept V3. In comparison, existing comparison methods experience severe catastrophic forgetting, resulting in subpar videos with distorted motions and inconsistent identity representation.
Figure 14: Comparison results of multi-concept video customization under the CTVC setting when the backbone is DreamVideo.
Multi-Concept Video Customization: Fig. 14 illustrates detailed comparative experiments on multi-concept video customization. To ensure fairness in evaluating against existing approaches, we incorporate the region-aware cross-attention mechanism from Mix-of-Show (gu2023mixofshow) into the spatial transformer block (STB) of all SOTA competing methods. The qualitative analysis in Fig. 14 highlights the exceptional capability of our model in multi-concept video customization, especially when continuously acquiring new personalized concepts under the CTVC setting. Specifically, when provided with an initial text prompt and several region text prompts accompanied by user-defined bounding boxes, all competing methods display substantial concept neglect and struggle to achieve effective multi-concept video customization. Firstly, it is clear that these methods fail to retain the identity of previously learned concepts due to catastrophic forgetting. Secondly, they often overlook crucial subject concepts and motion patterns during the video generation process. For instance, while users might aim to synthesize a video featuring V5 and V10 performing V26 together, none of the existing comparative methods successfully generate both V10 and V26, as depicted in Fig. 14.
Figure 15: Comparison results of video style transfer under the CTVC setting when the backbone is DreamVideo.
Style Transfer: As illustrated in Fig. 15, we present comparative analyses of customized video style transfer to assess the efficacy of our proposed CCVD in tackling the challenges of the CTVC problem. From the results in Fig. 15, it is evident that our proposed model outperforms other existing methods in generating videos that align with user-provided style concepts. This demonstrates the critical role of the designed task-aware concept aggregation module in maintaining unique style attributes under the CTVC setting. In contrast, other competing methods face difficulties in capturing the identity of diverse style concepts, primarily due to the challenges of catastrophic forgetting and concept neglect in the CTVC setting.
Societal Impact and Limitations
Societal Impact: Continual text-to-video customization (CTVC) has the potential to transform content creation by making high-quality personalized video production more accessible. This technology benefits various industries, including education, entertainment, marketing, and training, by allowing users to generate videos tailored to their specific needs. For educators, it enables the dynamic creation of instructional materials that adapt to different learning styles, making education more engaging and personalized. In entertainment, independent creators and small businesses can leverage this tool to produce unique, professional-grade videos without requiring expensive equipment or specialized expertise. Furthermore, continual learning capabilities allow models to refine and expand their knowledge over time. Users can introduce new subjects, motion patterns, or artistic styles while maintaining previously learned concepts, resulting in a more flexible and adaptive video generation process. This personalization enhances user experience, enabling iterative improvements based on real-time feedback.
Limitations: While CTVC offers numerous advantages, it also has some limitations. One is copyright infringement, as users might finetune models on proprietary materials without authorization. Another is privacy concerns, as models may require storing user preferences and personal data for customization. Balancing flexibility and privacy protection is crucial for the long-term success of this technology.